OpenStudy (chasebraves):

Simplify the expression where possible 1/(x^4)^-2

OpenStudy (trojanpoem):

1/(x^4)^-2 ( you multiply the power) 1/x^-8 (negative power = 1/) 1/1/x^8 = x^8

OpenStudy (chasebraves):

thanks

OpenStudy (trojanpoem):

any time.